OpenStudy (anonymous):

solve cubed root of 2 times the square root of 3 times the 4th root of two

OpenStudy (whpalmer4):

\[\sqrt[3]{2}*\sqrt{3}*\sqrt[4]{2}\]?

OpenStudy (anonymous):

yes @whpalmer4

OpenStudy (whpalmer4):

There isn't anything to solve there, are we supposed to simplify it?

OpenStudy (anonymous):

yeah, sorry thats what i meant

OpenStudy (whpalmer4):

Rewrite it in exponential form: \[\sqrt[n]{x} = x^{1/n}\] Then recall that \[a^n*a^m = a^{n+m}\]

OpenStudy (anonymous):

okay i wrote them in exponential form but i dont get it after that

OpenStudy (whpalmer4):

What do you have for your exponential form?

OpenStudy (anonymous):

2^(1/3) * 3^(1/2) * 2^(1/4)....is that right?

OpenStudy (whpalmer4):

\[2^{1/3}*3^{1/2}*2^{1/4}\]I see two parts of that with a common base, remember that \[a^n*a^m = a^{n+m}\]